# Wasserstein–Aitchison GAN for angular measures of multivariate extremes

### v2

@@ 4.2 Simulated data
As an illustration of the method’s performance in a controlled setting, we begin by considering simulated data. In Section 4.2.1, we adopt the same experimental setup as in Section 3.2 of [Girard-et-al], with which we compare our method. The dependence structure is logistic, characterized by a single parameter that controls the strength of dependence. This scenario is particularly simple in the sense that each sub-vector exhibits the same form of dependence, and increasing the dimension does not necessarily make the problem more difficult, since the algorithm can easily learn from this highly symmetric structure. For this reason, we consider in Section 4.2.2 a Hüsler–Reiss dependence structure, where each of the d(d1)/2d(d-1)/2 pairs of random variables in $\bm{X}$ may have a different level of dependence. Finally, in Section 4.2.3, we turn to an asymptotically independent copula—the Gaussian copula—as the dependence structure of $\bm{X}$. In this case, Assumption 2 is no longer satisfied, since the angular measure $\Phi$ is concentrated on the vertices of the simplex. Here, our interest lies in assessing whether our methodology can be improved by allowing for a heavier-tailed latent distribution in the generator.
@@ 4.2.1 Logistic dependence structure
We consider various scenarios in which we assess the performance of WA-GAN and compare it to HTGAN and GPGAN. For HTGAN we follow [Girard-et-al] and make the unrealistic assumption that $\alpha$ is known. We take $d\in\{10,20,50\}$ and $\theta\in\{4/3,2,4\}$ leading to a Kendall’s tau of $\tau\in\{1/4,1/2,3/4\}$. For the margins, we consider $\alpha=2$ in any scenario. The models are trained on $n_{\text{train}}=10\,000$ observations and validated on $n_{\text{val}}=5\,000$ observations. The performance assessment is done on $n_{\text{test}}=20\,000$ points.
@@ Results.
Some additional comments can be made for each method. For WA-GAN, the penalization coefficient $\rho$ in (19) is typically selected to be large among the possible values ($\rho=1$ or $\rho=3$) showing that it is a good idea to enforce marginal constraints (3) in the generative model. For HTGAN, the obtained values of the hyperparameters match the discussion on pages 13–15 in the paper [Girard-et-al] proposing this method. In particular, we get big batch sizes, big latent space dimensions, and low values of the learning rate included in the range recommendend by the authors. For GPGAN, the training phase is quicker since, as illustrated on Figure 2, we have a lower training size in this case as this method is trained on points with $|\bm{V}|_{\infty}>n/k$ while WA-GAN is trained on points with $|\bm{V}|_{1}>n/k$, which is less restrictive. This can be an asset if we have enough data, but can harm GPGAN’s performance if $d$ is large and $n$ is not too big.
As an illustration of the performances of the different methods to accurately reproduce the tail dependence of the data, we plot the empirical bivariate and trivariate extremal coefficients ($|J|=2$ and $|J|=3$ in (21)) of the test set against the ones of the training set, and against those of all the generative methods for $d\in\{10,50\}$ and $\tau=0.5$. In this setting, we know that the true values of these coefficients are $\sqrt{|J|}$. This is displayed n Figure 3 for $d\in\{10,50\}$. There is no clear visual difference between WA-GAN and HTGAN, both performing well at representing the coefficients of the test set. GPGAN, however, fails to do so, and even more when the dimension increases. When $d$ gets larger, the empirical extremal coefficients are below their true values, even in the test set. The reason is that the empirical estimator of $\Phi$ in (21) becomes less accurate as $d$ increases [clemencon2023concentration].
@@
Next, to illustrate the accuracy of generated extremes by each method, we plot the projection of each generated sample on the first two margins and compare it to the first two margins of the extremes in the test set. We take again $\tau=0.5$. Results for are displayed on Figure 4. WA-GAN seems to perform well, whatever the dimension. The results of HTGAN are less satisfying as it is clear from the figure that the angular measure associated with this method is discrete (this corresponds to the “directions” that we observe in the generated extremes) and this does not match the true angular measure which is logistic. This is theoretically justified by Proposition 2.10 in [Girard-et-al] introducing this method, and already visible from their Figure 2 on page 14. GPGAN also performs well even though its performance decreases in comparison to WA-GAN as the dimension increases; see Table 1.
@@ 4.2.2 Hüsler–Reiss dependence structure
In this setup, the random vector $\bm{X}$ is generated using a $d$-dimensional multivariate Pareto distribution with Hüsler–Reiss dependence structure having a random parameter matrix and Pareto(2) margins. The parameter matrix generation and the random sampling are done using the R package graphicalExtremes [graphicalExtremes]. The family is parametrized by a variogram matrix $\Gamma=(\gamma_{ij})_{i,j=1}^{d}$ where $\gamma_{ij}=0$ means complete dependence and $\gamma_{ij}=+\infty$ means asymptotic independence.
@@ 4.3 Financial data: daily returns of industry portfolios
We apply our methodology to analyze the tail behavior of the “value-averaged” daily returns of $d=30$ industry portfolios compiled and posted as part of the Kenneth French Data Library. The data in consideration span between 1950 and 2015 with n=16 694n=16\,694 observations. This data set is very widely used; for analyses in an extreme value context see e.g. [janssen2020] and [cooley2019]. Details about the portfolio constructions can be found online. [^1] https://mba.tuck.dartmouth.edu/pages/faculty/ken.french/Data_Library/det_30_ind_port.html Since we are interested in extreme losses we first multiply all returns by 1-1.
@@
To sample extremes from the joint distribution, we first compute the maximum likelihood estimators of the univariate GP parameters for each margin. Boxplots illustrating the values of those parameters are reported in Figure 12 in Appendix E. The estimated shape values clearly indicates the presence of heavy tails in the marginal distributions. The GP qq-plots of the marginal fits are to be found on Figures 13, 14 and 15 in Appendix E in the supplement. They indicate a relatively good fit, even though some extreme quantiles are larger than the ones associated with the GP model.
@@
As already pointed out in the analysis of [janssen2020], the extreme losses of those portfolios can be clustered into different kind of industries. For example, the tobacco industry exhibits asymptotic independence to all other categories (e.g., to the textile industries) while the extreme losses of the business and IT-related industries are dependent. We illustrate the fact that WA-GAN also captures those phenomena in Figure 10(a) by displaying the two-dimensional projections of the estimated angular measure on those margins. We see that the generated “angles” associated with tobacco/textile portfolios are much more concentrated around the axes than for business/IT, showing that WA-GAN is able to capture quite reasonably this complex tail dependence in this setting.
As a further illustration we consider a portfolio obtained by combining the mines industry portfolio with the coal industry portfolio (with equal weight). It is intuitive and has been shown in [cooley2019, janssen2020] that the mines and coal industries exhibit asymptotic dependence. If one wants to sample the losses associated to the portfolio we propose, this tail dependence should be taken into account. Figure 10(b) represents the densities of simulated values of the extreme losses of the combined portfolio (i.e., both marginal losses exceeds a large threshold) based on WA-GAN (in black) and based on independent simulations from the marginal GP distributions (in blue). Not taking the tail dependence into account leads to underestimation of the risk associated with this portfolio.
@@ 5 Conclusions
Another important aspect, not addressed in this work but potentially relevant in applications, is the possible time-evolving nature of the extremal dependence structure of the data [drees2023time, deCarvalho2025techs], or its variation with respect to a latent variable. Covariates can be incorporated into (Wasserstein) GANs by using their conditional versions [mirza2014cGAN, zheng2020cWGANGP], which could facilitate the generation of new multivariate extremes for specific covariate values, like time, in a manner similar to what has been explored for image generation in [gauthiercGAN]. These directions are left for future work.
@@ Appendix A Proof of Proposition 1
The exponent measure $\nu$ is connected to the angular measure $\Phi$ introduced in (2). More precisely [beirlant2004, Chapter 8], for any bounded, measurable function $f:\mathbb{E}\rightarrow\mathbb{R}$ that is zero in a neighborhood of the origin, we have
@@
From the local uniformity of the convergence in point (ii) of Remark 1, guaranteed by Assumption 3, and the fact that $\mathsf{P}(Y_{j}>0)=1$ for any $j\in\{1,\ldots,d\}$ by the previous computations and Assumption 2, the extended continuous mapping theorem [van1996weak, Theorem 1.11.1] applies, yielding
@@ Appendix B The Aitchison simplex
The open simplex ${\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1}$ is equipped with an inner product space structure with the following operations [aitchison1982]: for $\bm{v}=(v_{1},\ldots,v_{d}),\bm{w}=(w_{1},\ldots,w_{d})\in{\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1}$ and $\alpha\in\mathbb{R}$, define the following operations:
@@ Appendix C Angular measure with respect to another norm
for any Borel set $A\subseteq\widetilde{\Delta}_{d-1}$ such that $\widetilde{\Phi}(\partial A)=0$. Popular choices for $\|\cdot\|$ consist of the $L_{p}$ norms $|\cdot|_{p}$ on $\mathbb{R}^{d}$ for $p\in[1,\infty]$, see, e.g., [einmahl2001nonparametric, einmahl2009maximum, clemencon2023concentration]. Even though we only consider the norm $|\cdot|_{1}$ in Algorithm 1, we show that a simple post-processing step permits to provide an estimate for $\widetilde{\Phi}$ also.
@@
Appendix D Implementation of competing methods
HTGAN is follows a standard GAN architecture [goodfellow2014] with heavy-tailed latent variables that are Pareto($\alpha$) distributed for some $\alpha>0$. It does not make use of (multivariate) extreme value theory at all and rests on the capacity of the GAN to accurately capture the dependence of the data, also in the tail. As advised in their paper, we base our implementation on the one in [github_GAN]. Their preprocessing step in their Algorithm 1 on page 10, equation (3.11), requires an estimate of $\gamma=1/\alpha$ which is not provided in their methodology. They use the true value in their simulations, which is why we will also do so here, in Section 4.2. For real data applications, we will consider $\gamma=1/1.5$ as recommended by the authors in their results on page 19. Note also that it is not clear to us how step (3.12) in their Algorithm 2 is supposed to work as $\tilde{X}_{j}^{(i)}$ is not an element of $[0,1]$ so that it does not make sense to apply the inverse of the estimated marginal distribution function $F_{j}$. Since in this simulation setting the output of their generator is supposed to live in the same space as the original data, we do not apply any transform to compute the score (23). The score (22) is, of course, not affected by any marginal transformation. In real data applications, we do not compute the $W_{2}$ score and only consider the dependence score as done in the original paper. In terms of architectures and hyperparameters, we consider the same setting as for our method. This is very close to what the authors also do in their paper, besides the hyperparameter search which is performed through a Bayesian search in their numerical section. Both the generator and the discriminator are trained on $5000$ epochs. The final layer of the discriminator has a sigmoid activation function, as it is commonly applied in GAN architectures.
The GPGAN is closer to our approach as it is also based on the MGP distribution to accurately sample from the tail of $\bm{X}$, even though the approach in [Li-et-al] is quite different. The authors also rely on a GAN architecture so that we also rely on [github_GAN] to implement their methodology. The main differences between their approach and ours are as follows:
  • In [Li-et-al], the dependence is captured by the spectral vector $\bm{S}$ as introduced in (11), while our techniques rely on sampling from $\Phi$ after a transformation to the Aitchison coordinates which helps simplifying the distributions.
  • The discriminator in the GPGAN procedure directly evaluates sample quality on the MGP scale (after evaluating the marginal parameters), while our technique separates the dependence structure, for which we rely on the WGAN architecture (Algorithm 1), and the marginal considerations that permit to transform the output back to the MGP scale (Algorithm 2). Our approach has the advantage that if one is only interested in evaluating a tail dependence measure, like the extremal coefficients in (22), it is possible to do so without putting any assumptions on the marginal tail behavior and to work only with Algorithm 1.
The adaptive procedure in [Li-et-al] to pick the threshold $\bm{u}$ based on the “extremeness” function $E$ was not entirely clear to us. Therefore, we decided to pick the same $\bm{u}$ as we do to compute the marginal parameters, so that we work with the same estimates of the parameters $\bm{\sigma}$ and $\bm{\xi}$. Also, as in [Li-et-al] the latent distribution is not indicated, we consider a multivariate standard normal distribution on the latent space. The architectures and hyperparameters are also treated in the same way as for our method and the HTGAN method. Both the generator and the discriminator are trained on $5000$ epochs, based on the Adam optimizer, as done in the original paper [Li-et-al]. The final layer of the discriminator has a sigmoid activation function, as it is commonly applied in GAN architectures. Note that computing the score (23) poses no problem with this method as it outputs directly on the scale of $\bm{X}\mid\bm{X}\nleq\bm{u}(t)$; however, for the score (22), one needs to be more cautious. For this, we rely on the fact that if $\bm{X}^{*}_{1},\ldots,\bm{X}^{*}_{n}$ forms a sample from $\bm{X}\mid\bm{X}\nleq\bm{u}(t)$, then
$\bm{V}^{*}_{i}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\bm{b}(\bm{X}^{*}_{i})\sim\bm{V}\mid\bm{V}\nleq t=\bm{V}\mid|\bm{V}|_{\infty}>t,\qquad i\in\{1,\ldots,n\}.$
For $t>1$ large enough, it means that
$\bm{\Theta}_{i}^{*}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{\bm{V}^{*}_{i}}{|\bm{V}^{*}_{i}|_{\infty}}\stackrel{{\scriptstyle\mathcal{L}}}{{\approx}}\Phi_{\infty},\qquad i\in\{1,\ldots,n\}$
where $\Phi_{\infty}$ the angular measure with respect to the $L_{\infty}$-norm. Relying on the change of norm explained in Appendix C, we get an estimate for $\Phi$ by considering the discrete measure
$\widehat{\Phi}^{*}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{i=1}^{n}\Lambda_{i}^{*}\delta_{\frac{\bm{\Theta}_{i}^{*}}{|\bm{\Theta}_{i}^{*}|_{1}}},\qquad\Lambda_{i}^{*}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{|\bm{\Theta}_{i}^{*}|_{1}}{\sum_{i=1}^{n}|\bm{\Theta}_{i}^{*}|_{1}}.$
Appendix E Additional figures
(a) $d=10$
(b) $d=20$
(c) $d=50$
Figure 11: Heatmaps of the variogram matrices used in the simulations of Section 4.2.2.
Figure 12: Boxplots of estimated marginal GP parameters
Figure 13: GP qq-plots for margins 1 to 10 in the financial dataset.
Figure 14: GP qq-plots for margins $1$ to 20 in the financial dataset.
Figure 15: GP qq-plots for margins 21 to 30 in the financial dataset.

### v3

@@ 4.2 Simulated data
As an illustration of the method’s performance in a controlled setting, we begin by considering simulated data. In Section 4.2.1, we adopt the same experimental setup as in Section 3.2 of [25], with which we compare our method. The dependence structure is logistic, characterized by a single parameter that controls the strength of dependence. This scenario is particularly simple in the sense that each sub-vector exhibits the same form of dependence, and increasing the dimension does not necessarily make the problem more difficult, since the algorithm can easily learn from this highly symmetric structure. For this reason, we consider in Section 4.2.2 a Hüsler–Reiss dependence structure, where each of the d(d1)/2d(d-1)/2 pairs of random variables in $\bm{X}$ may have a different level of dependence. Finally, in Section 4.2.3, we turn to an asymptotically independent copula—the Gaussian copula—as the dependence structure of $\bm{X}$. In this case, Assumption 2 is no longer satisfied, since the angular measure $\Phi$ is concentrated on the vertices of the simplex. Here, our interest lies in assessing whether our methodology can be improved by allowing for a heavier-tailed latent distribution in the generator.
@@ 4.2.1 Logistic dependence structure
We consider various scenarios in which we assess the performance of WA-GAN and compare it to HTGAN and GPGAN. For HTGAN we follow [25] and make the unrealistic assumption that $\alpha$ is known. We take $d\in\{10,20,50\}$ and $\theta\in\{4/3,2,4\}$ leading to a Kendall’s tau of $\tau\in\{1/4,1/2,3/4\}$. For the margins, we consider $\alpha=2$ in any scenario. The models are trained on $n_{\text{train}}=10\,000$ observations and validated on $n_{\text{val}}=5\,000$ observations. The performance assessment is done on $n_{\text{test}}=20\,000$ points.
@@ Results.
Some additional comments can be made for each method. For WA-GAN, the penalization coefficient $\rho$ in (19) is typically selected to be large among the possible values ($\rho=1$ or $\rho=3$) showing that it is a good idea to enforce marginal constraints (3) in the generative model. For HTGAN, the obtained values of the hyperparameters match the discussion on pages 13–15 in the paper [25] proposing this method. In particular, we get big batch sizes, big latent space dimensions, and low values of the learning rate included in the range recommendend by the authors. For GPGAN, the training phase is quicker since, as illustrated on Figure 2, we have a lower training size in this case as this method is trained on points with $|\bm{V}|_{\infty}>n/k$ while WA-GAN is trained on points with $|\bm{V}|_{1}>n/k$, which is less restrictive. This can be an asset if we have enough data, but can harm GPGAN’s performance if $d$ is large and $n$ is not too big.
As an illustration of the performances of the different methods to accurately reproduce the tail dependence of the data, we plot the empirical bivariate and trivariate extremal coefficients ($|J|=2$ and $|J|=3$ in (21)) of the test set against the ones of the training set, and against those of all the generative methods for $d\in\{10,50\}$ and $\tau=0.5$. In this setting, we know that the true values of these coefficients are $\sqrt{|J|}$. This is displayed n Figure 3 for $d\in\{10,50\}$. There is no clear visual difference between WA-GAN and HTGAN, both performing well at representing the coefficients of the test set. GPGAN, however, fails to do so, and even more when the dimension increases. When $d$ gets larger, the empirical extremal coefficients are below their true values, even in the test set. The reason is that the empirical estimator of $\Phi$ in (21) becomes less accurate as $d$ increases [12].
@@
Next, to illustrate the accuracy of generated extremes by each method, we plot the projection of each generated sample on the first two margins and compare it to the first two margins of the extremes in the test set. We take again $\tau=0.5$. Results for are displayed on Figure 4. WA-GAN seems to perform well, whatever the dimension. The results of HTGAN are less satisfying as it is clear from the figure that the angular measure associated with this method is discrete (this corresponds to the “directions” that we observe in the generated extremes) and this does not match the true angular measure which is logistic. This is theoretically justified by Proposition 2.10 in [25] introducing this method, and already visible from their Figure 2 on page 14. GPGAN also performs well even though its performance decreases in comparison to WA-GAN as the dimension increases; see Table 1.
@@ 4.2.2 Hüsler–Reiss dependence structure
In this setup, the random vector $\bm{X}$ is generated using a $d$-dimensional multivariate Pareto distribution with Hüsler–Reiss dependence structure having a random parameter matrix and Pareto(2) margins. The parameter matrix generation and the random sampling are done using the R package graphicalExtremes [22]. The family is parametrized by a variogram matrix $\Gamma=(\gamma_{ij})_{i,j=1}^{d}$ where $\gamma_{ij}=0$ means complete dependence and $\gamma_{ij}=+\infty$ means asymptotic independence.
@@ 4.3 Financial data: daily returns of industry portfolios
We apply our methodology to analyze the tail behavior of the “value-averaged” daily returns of $d=30$ industry portfolios compiled and posted as part of the Kenneth French Data Library. The data in consideration span between 1950 and 2015 with n=16 694n=16\,694 observations. This data set is very widely used; for analyses in an extreme value context see e.g. [34] and [13]. Details about the portfolio constructions can be found online. [^1] https://mba.tuck.dartmouth.edu/pages/faculty/ken.french/Data_Library/det_30_ind_port.html Since we are interested in extreme losses we first multiply all returns by 1-1.
@@
To sample extremes from the joint distribution, we first compute the maximum likelihood estimators of the univariate GP parameters for each margin. Boxplots illustrating the values of those parameters are reported in Figure 2 in Appendix E. The estimated shape values clearly indicates the presence of heavy tails in the marginal distributions. The GP qq-plots of the marginal fits are to be found on Figures 3, 4 and 5 in Appendix E in the supplement. They indicate a relatively good fit, even though some extreme quantiles are larger than the ones associated with the GP model.
@@
As already pointed out in the analysis of [34], the extreme losses of those portfolios can be clustered into different kind of industries. For example, the tobacco industry exhibits asymptotic independence to all other categories (e.g., to the textile industries) while the extreme losses of the business and IT-related industries are dependent. We illustrate the fact that WA-GAN also captures those phenomena in Figure 10(a) by displaying the two-dimensional projections of the estimated angular measure on those margins. We see that the generated “angles” associated with tobacco/textile portfolios are much more concentrated around the axes than for business/IT, showing that WA-GAN is able to capture quite reasonably this complex tail dependence in this setting.
As a further illustration we consider a portfolio obtained by combining the mines industry portfolio with the coal industry portfolio (with equal weight). It is intuitive and has been shown in [13, 34] that the mines and coal industries exhibit asymptotic dependence. If one wants to sample the losses associated to the portfolio we propose, this tail dependence should be taken into account. Figure 10(b) represents the densities of simulated values of the extreme losses of the combined portfolio (i.e., both marginal losses exceeds a large threshold) based on WA-GAN (in black) and based on independent simulations from the marginal GP distributions (in blue). Not taking the tail dependence into account leads to underestimation of the risk associated with this portfolio.
@@ 5 Conclusions
Another important aspect, not addressed in this work but potentially relevant in applications, is the possible time-evolving nature of the extremal dependence structure of the data [14, 17], or its variation with respect to a latent variable. Covariates can be incorporated into (Wasserstein) GANs by using their conditional versions [40, 58], which could facilitate the generation of new multivariate extremes for specific covariate values, like time, in a manner similar to what has been explored for image generation in [24]. These directions are left for future work.
@@ Acknowledgments
The authors sincerely thank the anonymous reviewers for their insightful comments and suggestions, which have significantly improved the quality and clarity of this paper.
@@ Appendix A Proof of Proposition 1
The exponent measure $\nu$ is connected to the angular measure $\Phi$ introduced in (2). More precisely [7, Chapter 8], for any bounded, measurable function $f:\mathbb{E}\rightarrow\mathbb{R}$ that is zero in a neighborhood of the origin, we have
@@
From the local uniformity of the convergence in point (ii) of Remark 1, guaranteed by Assumption 3, and the fact that $\mathsf{P}(Y_{j}>0)=1$ for any $j\in\{1,\ldots,d\}$ by the previous computations and Assumption 2, the extended continuous mapping theorem [55, Theorem 1.11.1] applies, yielding
@@ Appendix B The Aitchison simplex
The open simplex ${\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1}$ is equipped with an inner product space structure with the following operations [1]: for $\bm{v}=(v_{1},\ldots,v_{d}),\bm{w}=(w_{1},\ldots,w_{d})\in{\kern 0.0pt\Delta}^{\mathrm{o}}_{d-1}$ and $\alpha\in\mathbb{R}$, define the following operations:
@@ Appendix C Angular measure with respect to another norm
for any Borel set $A\subseteq\widetilde{\Delta}_{d-1}$ such that $\widetilde{\Phi}(\partial A)=0$. Popular choices for $\|\cdot\|$ consist of the $L_{p}$ norms $|\cdot|_{p}$ on $\mathbb{R}^{d}$ for $p\in[1,\infty]$, see, e.g., [19, 21, 12]. Even though we only consider the norm $|\cdot|_{1}$ in Algorithm 1, we show that a simple post-processing step permits to provide an estimate for $\widetilde{\Phi}$ also.